Here is a calculation which is false but we have to find where is the mistake take place......Plz tell me that mistake with a satisfied reason....


 


 


Let....


 


2x= x + x


 


3x= x + x + x


 


nx= x + x + x +..........+x (n.times)


 


 


there fore


 


xx= x + x + x +............+x (x.times)


 


x2=x + x + x +........+ x


 


diff w.r.t. x both side


 


d(x2)/dx= d(x + x + x +........+ x)/dx


 


2x= 1+1+1+..........+1


 


2x=1*x   bcoz it is x times in above 3 equations


 


2=1


 


 


 


plzzzzzzz sir tell me the mistake of this qus with reason,

Grade:12

3 Answers

Tushar Makkar
12 years ago

n was a no. multiplied with x but when x is multiplied by x this thing is different as x is a variable and not constant.

u r  differentiating w.r.t x  ....till 2x=1+1+...+1 is correct but u can not write 1+1+1+.....x times =x as x is variable not a constant( u can write 1+1+1+.... n times = n because n is not a constant)it is hypothetically correct but mathematically incorrect... by doing this a variable will be introduced from a constant thing means 1+1+1.... is constant but x is variable

 

pls. approve the ans. if u agree with my opinion

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

There is error in differentiating RHS...

sheik arbaz
11 years ago

Hi friend,

             that is exactly correct.Your brialliant

             please,please,approave it.